OpenStudy (anonymous):

1/11in 1/2in 1/5in LCD :110? what do i do next?

OpenStudy (anonymous):

nexr step diya not yessssss

OpenStudy (turingtest):

multiply each fraction by (x/x) such that all denominators become the same: 1/11(10/10)=10/110 1/2(55/55)=55/110 1/5(22/22)=22/110

OpenStudy (anonymous):

hahaha..............

OpenStudy (anonymous):

\[|8-7\times5|+7\times \left| \left( -2^{3} \right) \right|\div4\]